OpenStudy (anonymous):

(sqrt x + 2 sqrt y)(sqrt x - 3 sqrt y)

OpenStudy (anonymous):

\[(\sqrt{x}+2\sqrt{y})(\sqrt{x}-3\sqrt{y})\]

OpenStudy (anonymous):

i was going to distribute